# JobFair: A Framework for Benchmarking Gender Hiring Bias in Large Language Models

### v1

@@ Appendix D Impact Ratio of Males Using Scoring Method with Mean As Cutoff
Appendix E Moving Average Comparing Male and Female Scores
@@
Appendix F Statistical Results: Testing Level and Spread Biases
@@
GPT-3.5 Healthcare $8.220$ $7.930$ $8.280$ $0.11600$ $0.26300$
GPT-3.5 Finance $8.490$ $8.070$ $8.390$ $0.09630$ $0.44600$
GPT-3.5 Construction $8.360$ $8.160$ $8.370$ $0.36000$ $0.53600$
GPT-4 Healthcare $8.320$ $8.170$ $8.380$ $0.13100$ $0.26300$
GPT-4 Finance $8.260$ $8.230$ $8.380$ $0.31800$ $0.35700$
GPT-4 Construction $8.240$ $8.220$ $8.300$ $0.68700$ $0.58500$
GPT-4o Healthcare $7.870$ $7.910$ $8.290$ $0.02610$ $0.16100$
GPT-4o Finance $7.780$ $7.940$ $8.240$ $0.10600$ $0.63300$
GPT-4o Construction $7.600$ $7.750$ $8.080$ $0.16800$ $0.55700$
Gemini-1.5-Pro Healthcare $7.800$ $8.010$ $8.010$ $1.00000$ $0.54500$
Gemini-1.5-Pro Finance $7.560$ $7.810$ $7.950$ $0.50700$ $0.40100$
Gemini-1.5-Pro Construction $7.940$ $8.100$ $8.240$ $0.47800$ $0.55900$
Gemini-1.5-Flash Healthcare $6.870$ $6.800$ $7.130$ $0.24300$ $0.38900$
Gemini-1.5-Flash Finance $6.610$ $6.740$ $7.130$ $0.15600$ $0.20600$
Gemini-1.5-Flash Construction $6.940$ $7.250$ $7.360$ $0.66800$ $0.67500$
Llama3-8b-Instruct Healthcare $8.540$ $8.530$ $8.610$ $0.58900$ $0.35300$
Llama3-8b-Instruct Finance $8.590$ $8.600$ $8.640$ $0.81300$ $0.41700$
Llama3-8b-Instruct Construction $8.430$ $8.470$ $8.500$ $0.88700$ $0.48000$
Llama3-70b-Instruct Healthcare $8.520$ $8.440$ $8.590$ $0.33600$ $0.21900$
Llama3-70b-Instruct Finance $8.640$ $8.690$ $8.710$ $0.92700$ $0.43900$
Llama3-70b-Instruct Construction $8.450$ $8.500$ $8.480$ $0.95500$ $0.59300$
Mistral-Large Healthcare $8.660$ $8.630$ $8.790$ $0.03530$ $0.09380$
Mistral-Large Finance $8.590$ $8.610$ $8.660$ $0.68100$ $0.34000$
Mistral-Large Construction $8.530$ $8.560$ $8.640$ $0.43200$ $0.37600$
Claude-3-Haiku Healthcare $8.110$ $7.910$ $8.270$ $0.05740$ $0.17900$
Claude-3-Haiku Finance $8.340$ $8.300$ $8.510$ $0.11700$ $0.28300$
Claude-3-Haiku Construction $8.130$ $8.140$ $8.090$ $0.84100$ $0.70400$
Claude-3-Sonnet Healthcare $8.410$ $8.320$ $8.470$ $0.36800$ $0.21000$
Claude-3-Sonnet Finance $8.300$ $8.340$ $8.450$ $0.49400$ $0.41800$
Claude-3-Sonnet Construction $8.290$ $8.290$ $8.320$ $0.91900$ $0.58800$
Notes: This table presents the average scores for neutral (Column 3), male (Column 4), and female resumes (Column 5) for the entire sample within each industry (Column 2) for each LLM (Column 1). Column 6 provides the p-value from a Permutation test, conducted with $100,000$ permutations, testing the null hypothesis that the scores are equal between male and female groups. Column 7 provides the p-value from another Permutation test, also with $100,000$ permutations, testing the null hypothesis that the variances are equal between male and female groups. We use a significance level of $0.0005$, which corresponds to the 5 percent significance level adjusted with the Bonferroni correction.
Appendix G Statistical Results: Testing Statistical and Taste-Based Bias
@@
Appendix H Statistical Results: Testing Industry-Effect on Bias Performance of LLMs
@@
Appendix I Moving Average for Home Distance and Last Year Working Status, with GPT-4o

### v2

@@ Appendix E Impact Ratio of Males Using Scoring Method with Mean As Cutoff
Appendix F Moving Average Comparing Male and Female Scores
@@
Appendix G Statistical Results with Ranking-After-Scoring Method: Testing Level and Spread Biases
@@
Appendix H Statistical Results with Scoring Method: Testing Level and Spread Biases
@@
GPT-3.5 H $8.220$ $7.930$ $8.280$ $0.00002$ $0.26300$
GPT-3.5 F $8.490$ $8.070$ $8.390$ $0.00000$ $0.44600$
GPT-3.5 C $8.360$ $8.160$ $8.370$ $0.02360$ $0.53600$
GPT-4 H $8.320$ $8.170$ $8.380$ $0.00110$ $0.26300$
GPT-4 F $8.260$ $8.230$ $8.380$ $0.00342$ $0.35700$
GPT-4 C $8.240$ $8.220$ $8.300$ $0.03180$ $0.58500$
GPT-4o H $7.870$ $7.910$ $8.290$ $0.00002$ $0.16100$
GPT-4o F $7.780$ $7.940$ $8.240$ $0.00000$ $0.63300$
GPT-4o C $7.600$ $7.750$ $8.080$ $0.00000$ $0.55700$
Gemini-1.5-Pro H $7.800$ $8.010$ $8.010$ $1$ $0.54500$
Gemini-1.5-Pro F $7.560$ $7.810$ $7.950$ $0.00377$ $0.40100$
Gemini-1.5-Pro C $7.940$ $8.100$ $8.240$ $0.03412$ $0.55900$
Gemini-1.5-Flash H $6.870$ $6.800$ $7.130$ $0.00005$ $0.38900$
Gemini-1.5-Flash F $6.610$ $6.740$ $7.130$ $0.00003$ $0.20600$
Gemini-1.5-Flash C $6.940$ $7.250$ $7.360$ $0.09359$ $0.67500$
Llama3-8b-Instruct H $8.540$ $8.530$ $8.610$ $0.01070$ $0.35300$
Llama3-8b-Instruct F $8.590$ $8.600$ $8.640$ $0.81300$ $0.41700$
Llama3-8b-Instruct C $8.430$ $8.470$ $8.500$ $0.88700$ $0.48000$
Llama3-70b-Instruct H $8.520$ $8.440$ $8.590$ $0.33600$ $0.21900$
Llama3-70b-Instruct F $8.640$ $8.690$ $8.710$ $0.04493$ $0.43900$
Llama3-70b-Instruct C $8.450$ $8.500$ $8.480$ $0.18103$ $0.59300$
Mistral-Large H $8.660$ $8.630$ $8.790$ $0.00003$ $0.09380$
Mistral-Large F $8.590$ $8.610$ $8.660$ $0.02459$ $0.34000$
Mistral-Large C $8.530$ $8.560$ $8.640$ $0.00416$ $0.37600$
Claude-3-Haiku H $8.110$ $7.910$ $8.270$ $0.00005$ $0.17900$
Claude-3-Haiku F $8.340$ $8.300$ $8.510$ $0.00000$ $0.28300$
Claude-3-Haiku C $8.130$ $8.140$ $8.090$ $0.31975$ $0.70400$
Claude-3-Sonnet H $8.410$ $8.320$ $8.470$ $0.01292$ $0.21000$
Claude-3-Sonnet F $8.300$ $8.340$ $8.450$ $0.00070$ $0.41800$
Claude-3-Sonnet C $8.290$ $8.290$ $8.320$ $0.25889$ $0.58800$
Notes: This table presents the average scores for neutral (Column 3), male (Column 4), and female resumes (Column 5) for the entire sample within each industry (Column 2) for each LLM (Column 1). Column 6 provides the p-value from a t-test, testing the null hypothesis that the scores are equal between male and female groups. Column 8 provides the p-value from a Permutation test, with $100,000$ permutations, testing the null hypothesis that the variances are equal between male and female groups. Columns 7 and 9 show the Anderson’s sharpened q-values, adjusting the p-values for the multi-testing issues, for Level and Spread tests respectively.
Appendix I Statistical Results: Testing Statistical and Taste-Based Bias
@@
Appendix J Statistical Results: Testing Industry-Effect on Bias Performance of LLMs
@@
Appendix K Moving Average for Home Distance and Last Year Working Status, with GPT-4o
@@
Appendix M Resume Example
Figure 16: The resume example for an applicant in Construction sector